OpenStudy (mpj4):

Question on Binomial theorem: How to find the specific term that contains b^(some specific exponent) in (a+b)^n? I don't know if I wrote it right. But that's what the problems generally look like.

OpenStudy (mpj4):

Here's one example. Find the term that contains y^8 in the binomial expansion of (2x+y^2)^6

OpenStudy (mathmate):

The coefficient of the \(r^{th}\) term of of the expansion of \((a+b)^n\) is \((^n_r)=\frac{n!}{(n-r)!r!}\). The \(r^{th}\) term is therefore \((^n_r)=\frac{n!}{(n-r)!r!}a^rb^{n-r}\) For the above example, \(a=2x, b=y^2, n=6, r=4\) (because \((y^2)^4=y^8\).

OpenStudy (mpj4):

OOOOHHHH How could I forget xD

OpenStudy (mpj4):

Thanks a bunch.

OpenStudy (mathmate):

You're welcome! :)